OpenStudy (anonymous):

The time required to finish a test in normally distributed with a mean of 80 minutes and a standard deviation of 15 minutes. What is the probability that a student chosen at random will finish the test in more than 95 minutes? A. 82% B. 2% C. 34% D. 16%

OpenStudy (anonymous):

Find the \(z\) score first. \[ z =\frac{x-\mu }{\sigma }=\frac{95-80}{15}=1 \]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=z+score+1

OpenStudy (anonymous):

It says the right tailed value (probability that is it above 95) is \(\approx 0.1587\) That makes \(16\%\) a very good contender here.

OpenStudy (anonymous):

thank you